APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday September 17, 2006 (MJD 53996) ---=== OBSERVING TEAM ===--- Swing: Elena Malanushenko Night: Shannon Watters, Dan Long ---=== OBSERVING PLAN ===--- 5(13||(3! (Science!) ---=== OBSERVING SUMMARY ===--- Marginally photometric conditions throughout the night. We started imaging Segue strip 1188 S until conditions deteriorated (mu=325 to mu=345.5, including ramps). We switched to SN strip 82 S for the rest of the night (lambda=126 to lambda=-122.7, including ramps). We started the spooling of the SN data. We had to disable the ODH sensor on the telescope fork because it was alarming in failure mode. The Oxygen level was okay. Dan rebooted the astrometric crate because ccd60 seemed 'dead' in the focus loop. The reboot appeared to clear up the problem (log). Elena found an ugly smear on the b1 spectro camera. It is covering at least 30 fibers (problems; Problem Report 7117). It's been like that since at least Friday night. Dan found the purge flow meter for the spectrograph was very low so he increased the flow. There is still an artifact on the cloud camera images that we think affects the measurement of the sky photometricity (problems). Run 6366 didn't show up in the imagingSummary section of the log. If we try a 'i4l' command in iop run 6366 doesn't appear. We had this same problem last night with Run 6362. ---=== OBSERVING LOG ===--- Afternoon Checkout: ------------------- It is clear, almost photometric, T=9.8C, wind up to 14 mph, humidity 45. All systems are checked and they work fine. Spectro focus has been checked and good for T=10C. Cartridge 3 plate 2441 has been used for spectro checkout and flat (frame 40680) has red value of NGOODFIBER=291. Night: -------- 01:10Z On-telescope imager bias taken under dark conditions. 01:30Z Enclosure off. 02:17Z Strip 1188 S, SEGUE, Run 6366. Skippy happy from the beginning. In focus from frame 13. Seeing ~1.6" (acceptable). 02:17Z mu=318 02:32Z Ramp started, mu=325, frame 48. 02:40Z Science started, mu=327, frame 62. 03:52Z Ramp ended, mu=345.25, frame 183. Non-photometric sky since ~03:45 with increased seeing so we ended the run. 03:57Z Rebooted the astrometric crate because ccd60 was just following the secondary on the focusLoop so Dan suspected the crate needed a nudge. It seemed to work... 04:00Z Strip 82 S, SN, Run 6367. Skippy happy from the beginning. In focus from frame 8. Seeing ~1.4". 04:02Z Ramp started, lambda=126, frame 8. 04:10Z Science started, lambda=128, frame 19. 11:16Z Science ended, lambda=-125, frame 732. 11:24Z Ramped ended, lambda=-122.75, frame 748. endNight cried tears of joy ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 6364 22:27Z 22:53Z 100 O 42.58 49.12 58 ignore 6365 01:10Z 01:27Z 100 O 83.53 87.78 42 bias 6367 04:02Z 11:25Z 82 S 126.29 -122.63 756 science ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- 6366 8 4 0 0.000 0.000 0.000 0.000 -0.00000 0 0 6366 57 4 170 4.200 0.161 -6.400 0.155 -0.01603 80 45 6366 108 4 63 3.600 0.167 -6.400 0.152 -0.01595 90 45 6366 160 4 63 3.500 0.156 -7.300 0.154 -0.01561 101 45 6367 35 4 119 -2.400 0.158 3.000 0.136 -0.00232 181 57 6367 86 4 75 -3.100 0.185 2.600 0.151 0.00058 181 57 6367 136 4 55 -3.400 0.136 2.400 0.152 0.00108 181 57 6367 188 4 60 -3.600 0.147 2.100 0.185 0.00119 181 57 6367 239 4 44 -3.500 0.147 1.700 0.129 0.00010 181 57 6367 291 4 45 -3.700 0.148 1.200 0.122 0.00089 181 57 6367 341 4 36 -3.800 0.201 0.800 0.131 0.00021 181 57 6367 392 4 44 -3.600 0.131 1.000 0.151 0.00051 181 57 6367 444 4 39 -3.600 0.121 0.900 0.134 0.00084 181 57 6367 495 4 33 -3.400 0.100 1.400 0.136 0.00068 181 57 6367 547 4 50 -3.700 0.139 1.200 0.118 -0.00038 181 57 6367 599 4 71 -3.200 0.148 1.000 0.105 -0.00064 181 57 6367 650 4 67 -3.600 0.188 1.000 0.142 -0.00031 181 57 6367 702 4 63 -3.400 0.175 1.100 0.146 -0.00096 181 57 ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ 6367 471 1 57 359 9 -0.851 0.242 0.862 0.210 -0.00290 6367 471 1 57 359 15 -0.583 0.215 0.712 0.246 -0.00239 6367 471 1 57 359 14 -0.090 0.174 1.175 0.200 -0.00395 6367 471 1 57 359 11 0.123 0.100 1.142 0.160 -0.00384 6367 471 1 57 359 7 0.416 0.197 1.330 0.236 -0.00447 6367 506 1 57 359 7 -1.066 0.143 -0.850 0.183 0.00286 6367 506 1 57 359 9 -0.700 0.176 -0.591 0.261 0.00199 6367 506 1 57 359 16 -0.495 0.166 -0.464 0.260 0.00156 6367 506 1 57 359 10 -0.055 0.268 -0.195 0.251 0.00065 6367 506 1 57 359 16 0.067 0.257 0.009 0.189 -0.00003 6367 522 1 57 359 8 -1.086 0.331 -0.093 0.264 0.00031 6367 522 1 57 359 14 -0.762 0.243 0.237 0.279 -0.00080 6367 522 1 57 359 11 -0.302 0.267 0.462 0.175 -0.00155 6367 522 1 57 359 10 -0.029 0.300 0.401 0.221 -0.00135 6367 522 1 57 359 9 0.152 0.292 0.658 0.261 -0.00221 6367 558 1 57 359 16 -0.965 0.138 0.022 0.198 -0.00007 6367 558 1 57 359 15 -0.600 0.183 -0.043 0.191 0.00014 6367 558 1 57 359 12 -0.190 0.235 0.266 0.205 -0.00089 6367 558 1 57 359 12 0.112 0.188 0.510 0.241 -0.00171 6367 558 1 57 359 13 0.316 0.134 0.263 0.128 -0.00088 6367 574 1 57 359 13 -0.968 0.161 -0.474 0.305 0.00160 6367 574 1 57 359 17 -0.539 0.201 -0.177 0.243 0.00060 6367 574 1 57 359 5 -0.232 0.183 -0.179 0.224 0.00060 6367 574 1 57 359 18 -0.043 0.239 -0.081 0.266 0.00027 6367 574 1 57 359 13 0.413 0.251 0.084 0.348 -0.00028 6367 608 1 57 359 22 -1.243 0.096 -0.183 0.180 0.00062 6367 608 1 57 359 15 -0.914 0.095 -0.207 0.266 0.00070 6367 608 1 57 359 13 -0.621 0.165 -0.126 0.284 0.00042 6367 608 1 57 359 12 -0.339 0.066 -0.087 0.244 0.00029 6367 608 1 57 359 10 -0.025 0.146 -0.018 0.306 0.00006 6367 626 1 57 359 8 -1.123 0.390 -0.049 0.261 0.00016 6367 626 1 57 359 12 -0.721 0.292 0.223 0.203 -0.00075 6367 626 1 57 359 16 -0.566 0.265 0.277 0.203 -0.00093 6367 626 1 57 359 12 -0.297 0.295 0.445 0.175 -0.00150 6367 626 1 57 359 12 0.013 0.358 0.393 0.237 -0.00132 6367 661 1 57 359 9 -1.723 0.191 0.751 0.260 -0.00252 6367 661 1 57 359 15 -1.137 0.287 1.001 0.148 -0.00336 6367 661 1 57 359 14 -0.888 0.267 1.123 0.190 -0.00378 6367 661 1 57 359 10 -0.676 0.399 1.239 0.132 -0.00417 6367 661 1 57 359 13 -0.637 0.201 1.624 0.186 -0.00546 6367 679 1 57 359 14 -1.042 0.305 0.268 0.203 -0.00090 6367 679 1 57 359 15 -0.826 0.275 0.373 0.260 -0.00126 6367 679 1 57 359 16 -0.273 0.287 0.403 0.277 -0.00136 6367 679 1 57 359 13 -0.040 0.358 0.705 0.159 -0.00237 6367 679 1 57 359 13 0.347 0.164 0.535 0.203 -0.00180 6367 713 1 57 359 22 -1.554 0.449 0.003 0.309 -0.00001 6367 713 1 57 359 15 -1.302 0.388 0.217 0.397 -0.00073 6367 713 1 57 359 20 -1.003 0.358 0.358 0.365 -0.00120 6367 713 1 57 359 16 -0.500 0.364 0.608 0.267 -0.00204 6367 713 1 57 359 21 -0.364 0.465 0.489 0.293 -0.00164 6367 729 1 57 359 11 -1.029 0.132 -0.233 0.221 0.00078 6367 729 1 57 359 22 -0.647 0.168 0.012 0.173 -0.00004 6367 729 1 57 359 19 -0.221 0.146 0.275 0.220 -0.00093 6367 729 1 57 359 16 0.077 0.179 0.415 0.235 -0.00139 6367 729 1 57 359 22 0.299 0.219 0.652 0.278 -0.00219 ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n):y QA Procedures Done (y/n):n/a UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 40679, plate -9999 ------- 22:30 40679 0.0 bias ----- sequence 40680, plate 2441 ------- 22:57 40680 10.0 flat 22:59 40681 2.0 arc 23:08 40682 2.1 arc 23:10 40683 2.1 arc ----- sequence 40684, plate -9999 ------- 23:27 40684 0.0 bias ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL Stays: JL ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 20:14Z 52 30 6 356 (NNW) 1015 - Clear 20:46Z 53 27 6 266 (W) 1034 - Clear 21:18Z 57 31 6 279 (W) 874 - Clear 21:50Z 56 31 9 269 (W) 748 - Clear 22:23Z 56 30 1 254 (WSW) 605 - Clear 22:53Z 53 24 11 272 (W) 550 - Clear 23:26Z 54 27 12 322 (NW) 648 - Clear 23:56Z 54 28 6 266 (W) 642 - Clear 00:27Z 52 30 14 295 (WNW) 659 - Clear 00:59Z 50 29 9 270 (W) 677 - Clear 01:31Z 49 29 10 269 (W) 620 - Clear 02:27Z 49 28 18 285 (WNW) 624 - Clear 02:58Z 49 34 11 268 (W) 771 - Clear 03:28Z 48 35 15 261 (W) 752 - Clear 04:00Z 49 34 12 262 (W) 798 - Clear 04:31Z 48 34 13 264 (W) 797 - Clear 05:01Z 48 35 9 268 (W) 806 - Clear 05:31Z 49 32 13 275 (W) 776 - Clear 06:04Z 49 33 9 286 (WNW) 820 - Clear 06:35Z 48 33 11 308 (NW) 823 - Clear 07:05Z 48 33 10 351 (NNW) 789 - Clear 07:35Z 47 33 11 6 (N) 794 - Clear 08:06Z 47 29 2 10 (N) 960 - Clear 08:36Z 46 23 13 0 (N) 823 - Clear 09:07Z 46 15 12 2 (N) 586 - Clear 09:37Z 47 13 18 18 (NNE) 443 - Clear 10:09Z 46 13 12 6 (N) 275 - Clear 10:42Z 45 13 22 24 (NNE) 289 - Clear 11:12Z 44 11 20 12 (NNE) 190 - Clear 11:42Z 43 11 9 21 (NNE) 187 - Clear 12:14Z 43 8 21 17 (NNE) 155 - Clear ---=== TELESCOPE STATUS ===--- 23:50Z Spectro forced to fill. 00:10Z Enclosure doors opened, fans and blowers on, louvers opened. 01:30Z Enclosure off. 12:00Z Enclosure on, doors closed, fans and blowers off, louvers closed. Telescope stowed at 30 degrees with cart 3 (plate 2441) mounted. Counterweights at 250. We topped the spectro dewars and left it in manual mode (off). Spectro dewar 12:23Z 67 lb, 23.1 psi Imager dewar 12:10Z 184 lb, 20.4 psi The ODH sensor on the telescope fork is disabled (unplugged) ---=== SOFTWARE USED ===--- IOP/SOP: v4_26_0 Watcher: v2_34_0 MCP: v6_3_2 TPM: tpm_v3_3_0 AstroDa: v15_12_1 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_19 SoS: v5_1_4 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 24.6840 24.6750 Axial B 26.2560 26.2540 Axial C 25.3370 25.3490 Trans D 13.5660 13.5760 Lateral E 9.9441 9.9568 Lateral F 11.5189 11.5316 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: 12.6 Alt.: 30.430295 MIGS TONIGHT NOMINAL Axial A 1.5530 1.5550 Axial B 1.1280 1.1290 Axial C 1.1880 1.1910 Trans D -0.0610 -0.0530 GALILS Commanded: 1613992. 1594402. 1543151. -10550. -5800. Actual: 1614053. 1594719. 1543082. -10535. -5822. SETMIR VALUES SecDesOrient: 1257.00 20.00 -10.00 -100.00 176.64 SecOrient: 1256.92 19.90 -10.06 -100.47 176.63 ---=== PROBLEMS IN DETAIL ===--- The number of good fibers for spectro cameras b1 and r1 differ by 30-50 fibers. Elena checked the fits files and found a smear in the center of b1. Details can be found in Problem Report 7117. There is still an artifact in the cloud camera images that changes 'strength'. When the artifact is most visible it seems to correlate with a higher sky standard deviation. We think it may be making photometric conditions look non-photometric. sp1> Coll_motor_A -3988 Coll_motor_B -3815 Coll_motor_C -2910 sp2> Coll_motor_A -43645 Coll_motor_B -44845 Coll_motor_C -42845 ----------------------------- sp1> caget tpm_TM_PMSSAvg tpm_TM_TelairAvg.VALA tpm_TM_PMSSAvg 12.8671 tpm_TM_TelairAvg.VALA 14.31 ---------------------------- Expected tight T=10C. --------------------------- IDL 5.5a> collimate,40682 % Compiled module: COLLIMATE. Reading /data/spectro/53996/sdR-b1-00040682.fit speclog_version: Command not found. Reading /data/spectro/53996/sdR-b1-00040683.fit speclog_version: Command not found. COLLIMATE: Min offset = -0.30 pix COLLIMATE: Max offset = 1.06 pix COLLIMATE: RMS across CCD = 0.33 pix COLLIMATE: Mean offset = 0.35 pix COLLIMATE: Camera b1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of -12.6 degrees for spectro-1 COLLIMATE: (if red is already in focus) Log file = Collimate-53996-b1-00040682.log PostScript file = Collimate-53996-b1-00040682.ps Reading /data/spectro/53996/sdR-b2-00040682.fit speclog_version: Command not found. Reading /data/spectro/53996/sdR-b2-00040683.fit speclog_version: Command not found. COLLIMATE: Min offset = -0.06 pix COLLIMATE: Max offset = 0.80 pix COLLIMATE: RMS across CCD = 0.26 pix COLLIMATE: Mean offset = 0.35 pix COLLIMATE: Camera b2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of -12.6 degrees for spectro-2 COLLIMATE: (if red is already in focus) Log file = Collimate-53996-b2-00040682.log PostScript file = Collimate-53996-b2-00040682.ps Reading /data/spectro/53996/sdR-r1-00040682.fit speclog_version: Command not found. Reading /data/spectro/53996/sdR-r1-00040683.fit speclog_version: Command not found. COLLIMATE: Min offset = -0.79 pix COLLIMATE: Max offset = 0.09 pix COLLIMATE: RMS across CCD = 0.27 pix COLLIMATE: Mean offset = -0.32 pix COLLIMATE: Camera r1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 2911 steps for spectro-1 COLLIMATE: Issue SOP command: "sp1; mechPiston 2911" Log file = Collimate-53996-r1-00040682.log PostScript file = Collimate-53996-r1-00040682.ps Reading /data/spectro/53996/sdR-r2-00040682.fit speclog_version: Command not found. Reading /data/spectro/53996/sdR-r2-00040683.fit speclog_version: Command not found. COLLIMATE: Min offset = -0.92 pix COLLIMATE: Max offset = -0.09 pix COLLIMATE: RMS across CCD = 0.23 pix COLLIMATE: Mean offset = -0.35 pix COLLIMATE: Camera r2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 3164 steps for spectro-2 COLLIMATE: Issue SOP command: "sp2; mechPiston 3164" Log file = Collimate-53996-r2-00040682.log PostScript file = Collimate-53996-r2-00040682.ps % Program caused arithmetic error: Floating illegal operand IDL 5.5a> ------------------- no move sp1, sp2.